# Class-Based Expurgation Attains Csiszár’s Expurgated Source-Channel Exponent

### v2

@@ I Introduction
[…]
@@ III-B Dual-Domain
Consider a given partitioning of the source sequences as described in Lemma 2. Then, there exists a codebook such that for any choice of parameters $\{\lambda_{ic},\rho_{ic}\}$ satisfying $\lambda_{ic}>0$ and $\rho_{ic}\geq 1$,
$\begin{split}&p_{e}\,\dot{\leq}\,\sum_{c=1}^{m}\sum_{i\in\Lambda_{c}}\exp\!\bigg{(}\!\!-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{ic}\big{)}+\big{(}\lambda_{ic}-\rho_{ic}\big{)}{R}_{i}\\ &\hskip 99.58464pt-t{E}_{\mathrm{s}}\big{(}\lambda_{ic},P_{V}\big{)}\Big{]}\bigg{)}.\end{split}$ (15)
@@
The function $\overline{E}_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$ is the concave hull on the interval [1,)[1$\infty$)[ 1 , ∞ ) of the function $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$.
@@
If $E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})$ is concave in $\rho$, or equivalently, if the distribution in $\mathcal{Q}_{m}$ that maximizes $E_{\mathrm{x}}^{\prime}(\rho,Q)$ is independent of $\rho$, then $\overline{E}_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})=E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})$, and hence
$E_{J,\mathrm{ex}}^{\mathrm{dual}}(t,\mathcal{Q}_{m})=\sup_{\rho\geq 1}\Big{\{}E_{\mathrm{x}}^{\prime}(\rho,\mathcal{Q}_{m})-tE_{\mathrm{s}}(\rho,P_{V})\Big{\}}.$ (20)
@@ III-C Two-Class Partition of the Source Sequences
$\mathcal{A}_{1}$ =$=\left\{\boldsymbol{v}:\big{|}\mathcal{T}^{k}(\hat{P}_{\boldsymbol{v}})\big{|}\leq e^{nR_{0}}\right\}$ (23)
$\mathcal{A}_{2}$ =$=\left\{\boldsymbol{v}:\big{|}\mathcal{T}^{k}(\hat{P}_{\boldsymbol{v}})\big{|}>e^{nR_{0}}\right\},$ (24)
@@ IV Single Class Coding
Assume that all codewords are drawn independently and uniformly from the type class $\mathcal{T}^{n}(Q)$, for some given $Q\in\mathcal{P}_{n}(\mathcal{X})$. Then, there exists a codebook such that

### v3

@@ I Introduction
We study the transmission of non-equiprobable messages from a discrete memoryless source (DMS) with distribution $P^{k}(\boldsymbol{v})=\prod_{i=1}^{k}P_{V}(v_{i})$, where $\boldsymbol{v}=(v_{1},...,v_{k})\in\mathcal{V}^{k}$ is the source message, and $\mathcal{V}$ is a finite discrete alphabet. The channel is a discrete memoryless channel (DMC) given by $W^{n}(\boldsymbol{y}|\boldsymbol{x})=\prod_{i=1}^{n}W(y_{i}|x_{i})$, $\boldsymbol{x}=(x_{1},\ldots,x_{n})\in\mathcal{X}^{n}$ and $\boldsymbol{y}=(y_{1},\ldots,y_{n})\in\mathcal{Y}^{n}$, where $\mathcal{X}$ and $\mathcal{Y}$ are discrete alphabets with cardinalities $|\mathcal{X}|$ and $|\mathcal{Y}|$, respectively. An encoder maps the length-$k$ source message $\boldsymbol{v}$ to a length-$n$ codeword $\boldsymbol{x}_{\boldsymbol{v}}$, which is then transmitted over the channel. We refer to $t\triangleq k/n$ as the transmission rate. Based on the channel output $\boldsymbol{y}$ the decoder guesses which source message was transmitted.
@@ III-B Dual-Domain
Consider a given partitioning of the source sequences as described in Lemma 2. Then, there exists a codebook such that for any choice of parameters $\{\lambda_{i},\rho_{i}\}$ satisfying $\lambda_{i}>0$ and $\rho_{i}\geq 1$,
$\begin{split}&p_{e}\,\dot{\leq}\,\sum_{c=1}^{m}\sum_{i\in\Lambda_{c}}\exp\!\bigg{(}\!\!-n\Big{[}{E}_{\mathrm{x}}^{\prime}\big{(}{Q}_{c},\rho_{i}\big{)}+\big{(}\lambda_{i}-\rho_{i}\big{)}{R}_{i}\\ &\hskip 99.58464pt-t{E}_{\mathrm{s}}\big{(}\lambda_{i},P_{V}\big{)}\Big{]}\bigg{)}.\end{split}$ (15)
@@
The function $\overline{E}_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$ is the concave hull on the interval $1,\infty)$ of the function $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$.
@@
If $E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$ is concave in $\rho$, or equivalently, if the distribution in $\mathcal{Q}_{m}$ that maximizes $E_{\mathrm{x}}^{\prime}(\rho,Q)$ is independent of $\rho$, then $\overline{E}_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)=E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)$, and hence
$E_{J,\mathrm{ex}}^{\mathrm{dual}}(t,\mathcal{Q}_{m})=\sup_{\rho\geq 1}\Big{\{}E_{\mathrm{x}}^{\prime}(\mathcal{Q}_{m},\rho)-tE_{\mathrm{s}}(\rho,P_{V})\Big{\}}.$ (20)
@@ III-C Two-Class Partition of the Source Sequences
$\mathcal{A}_{1}$ $=\left\{\boldsymbol{v}:\big{|}\mathcal{T}^{k}(\hat{P}_{\boldsymbol{v}})\big{|}\leq e^{nR_{0}}\right\}$ (23)
$\mathcal{A}_{2}$ $=\left\{\boldsymbol{v}:\big{|}\mathcal{T}^{k}(\hat{P}_{\boldsymbol{v}})\big{|}>e^{nR_{0}}\right\},$ (24)
@@ IV Single Class Coding
There exist single-class codes with constant composition ensembles that satisfies